### v1

# On Detecting Cherry-picking in News Coverage Using Large Language Models

Abstract
Cherry-picking refers to the deliberate selection of evidence or facts that favor a particular viewpoint while ignoring or distorting evidence that supports an opposing perspective. Manually identifying instances of cherry-picked statements in news stories can be challenging, particularly when the opposing viewpoint’s story is absent. This study introduces Cherry, an innovative approach for automatically detecting cherry-picked statements in news articles by finding missing important statements in the target news story. Cherry utilizes the analysis of news coverage from multiple sources to identify instances of cherry-picking. Our approach relies on language models that consider contextual information from other news sources to classify statements based on their importance to the event covered in the target news story. Furthermore, this research introduces a novel dataset specifically designed for cherry-picking detection, which was used to train and evaluate the performance of the models. Our best performing model achieves an F-1 score of about 89% in detecting important statements when tested on unseen set of news stories. Moreover, results show the importance incorporating external knowledge from alternative unbiased narratives when assessing a statement importance.
1 Introduction
The erosion of citizens’ trust in the media has reached a critical stage, posing significant harm to the foundations of democracy Dan et al. (2021); Druckman and Parkin (2005); Chiang and Knight (2011). Research shows evidence indicating a decline in media trust among citizens  Strömbäck et al. (2020). Media losing trust can be attributed to multiple factors, among which, two important factors are the perception of political bias in media reporting Lee (2010) and consumers’ exposure to misinformation and disinformation Lee et al. (2023); Hameleers et al. (2022).
Cherry-picking, a widely used but often imperceptible technique in news reporting, is a device that aids bias and disinformation Vincent (2019); Paul and Elder (2019). It is generally defined as a logical fallacy that involves deliberately selecting (and/or exaggerating) facts that bolster a certain argument and omitting (and/or under-emphasizing) facts that support the opposing viewpoint Best (2004). Common aliases for cherry-picking known in literature are one-sideness Paul and Elder (2008), argument by half-truth Sproule (2001), and case-making Lee (1945).
Cherry-picking occurs in a variety of domains, such as media narratives, science and politics Best (2004). Cherry-picking can be observed across scales as large as in selecting which events to cover, and as subtle as in choosing specific words for a particular story. Furthermore, cherry-picking can manifest in various types of data. For example, for facts staffed with numbers and stats, cherry-picking can occur when specific segments of a trend-line Asudeh et al. (2020) or particular statistics Best (2004) are chosen to bolster a particular claim Wu et al. (2017). For facts that are visual (e.g., videos), cherry-picking occurs when only certain frames are selectively presented.
The focus of this study is cherry-picking in statements within news articles, particularly when only certain facts are included while others are taken out. A multitude of media outlets, including prominent and highly regarded ones, exhibit biases in various topics within domains such as politics and science Baron (2006). These media outlets often exercise caution in curating the facts included in their stories covering controversial topics Paul and Elder (2019, 2008). This selection of facts is capable of swaying the public opinion on crucial matters, such as climate change, vaccination, or elections. It can mislead readers into believing that the partial truth presented in these stories represents the entire truth Fleming (1995).
We introduce a novel method for automatically detecting cherry-picking in news reports by leveraging context derived from the covered event. Our method employs language models, including BERT and Longformer, trained to assess a statement’s importance based on context. The models are trained on the novel Cherry dataset we curated for this purpose. The proposed method shows promising results in predicting a statement’s importance given external contextual information. Our top-performing model achieves about 89% in terms of F-1 score and about 90% in terms of accuracy. To the best of our knowledge, this is the first study that addresses the detection of cherry-picking in text using computational approach. The key contributions of this work are as follows: (a) The formulation and modeling of the problem from a computational perspective. (b) The novel end-to-end cherry-picking detection approach that incorporates contextual information. (c) Additionally, we introduce a new cherry-picking detection dataset. All artifacts related to this work can be found at our anonymous Github repository https://github.com/emnlp-cherry/Cherry.
2 Related Work
The research on media bias can be classified into three levels according to text granularity: word-level, sentence-level, and article-level Chen et al. (2020). Nevertheless, distinguishing between these three levels is not always straightforward, as they often intersect and influence each other to a certain degree. For instance, word-level bias is dependent on sentence or article context, whereas sentence-level bias relies on the broader article context.
•  Word-level bias. The choice of words in news articles can influence the audience’s opinions towards a particular event Hamborg (2020). Word-level bias encompasses two primary categories: framing bias and epistemological bias Recasens et al. (2013). The connotations associated with the same words can vary based on the specific context in which they are employed Spinde et al. (2021b).
•  Sentence-level bias. Numerous studies focus on detecting bias at sentence level, considering it as the fundamental level that can be combined with article-level bias Spinde et al. (2021a). The study in Hube and Fetahu (2018) introduced an automated method for identifying biased statements from Wikipedia using lexicon of bias words. In Lei et al. (2022), the authors studied the task of identifying biased sentences in news articles that serve to explain the overall bias present throughout the whole article. Focusing on media’s sentiment toward target entities, Fan et al. (2019) examined sentence-level bias along with article context.
•  Article-level bias. For identifying political ideology in news articles, Baly et al. (2020) constructed a dataset of news articles annotated with left, center, and right leanings, and built and evaluated machine learning models using the dataset to predict the articles’ ideological leaning.
AllSides.com, AdFontesMedia.com, and MBFC are widely utilized sources in media bias analysis, providing researchers with tools and ratings to evaluate the political bias and credibility of news sources. They employ diverse methods, such as community feedback, analysts with varying viewpoints, and a combination of objective measures and subjective analysis, to evaluate bias in news outlets. Literature addressing bias and misinformation detection heavily relies on these three sources for bias categorization in tasks such as source annotation on social media platforms Weld et al. (2021), comparing the diffusion of news from reliable and non-reliable sources Cinelli et al. (2020), benchmarking automatic media bias monitors Ribeiro et al. (2018), and creating corpora for automatic media bias detection Chen et al. (2020).
While previous media bias research primarily focuses on predisposition towards a particular viewpoint, such as ideological leanings and sentiment bias, there has been limited research on the intentional practice of cherry-picking in news articles, which is a tactic employed to selectively present information that supports a certain bias. Although existing studies have explored the impact of cherry-picking in various contexts, such as computational fact-checking Wu et al. (2017); Lin et al. (2021) and detecting cherry-picking in trendlines Asudeh et al. (2020), they have not specifically addressed cherry-picking in news articles. While Ehrhardt et al. (2021) introduced a method to predict news articles’ ideological leaning by examining the omission of entities, our approach distinguishes itself by focusing on detecting cherry-picked statements in news articles, instead of analyzing co-occurrences of entities across different articles.
Diverging from previous research, our study stands out as the first comprehensive examination of cherry-picking in news articles, specifically analyzing the importance of statements extracted from news articles and analyze whether these statements are underemphasized by the articles, thus constituting instances of cherry-picking. In light of this, we have undertaken the construction of a new cherry-picking dataset, providing a valuable resource for evaluating the prevalence and impact of cherry-picked statements in news articles. Moreover, we developed a cherry-picking detection pipeline to identify cherry-picked statements. Our contribution could shed light on the manipulation of information in news articles. It not only enhances our understanding of mechanisms behind cherry-picking but also contributes to the promotion of more objective and informed journalism.
3 Cherry-picking Facts in News Stories: Problem Definition and Modeling
Cherry-picking happens when a news outlet prioritizes less crucial statements that align with their biases to appear in their narrative, and/or neglects more important relevant statements, especially those against their biases (Paul and Elder, 2008; Vincent, 2019; Brown, 1963). To offer a more illustrative portrayal of cherry-picked statements, Table 1 provides a juxtaposition of three authentic news stories from distinctively biased outlets covering the same event. Semantically similar statements from the three articles are aligned by row. Empty cells in the table indicate that the corresponding statement is not mentioned by the respective news source. Bias categorization of media outlets in comparison is based on MediaBiasFactCheck.com (MBFC).
Table 1 illustrates cherry-picking in journalism by showcasing statements from authentic news articles by distinctively biased news outlets, all covering the same event: U.S. President Donald Trump’s attendance at the annual Davos economic forum. The news outlets included in Table 1 are reuters.com (center), newsmax.com (right-bias), and cnbc.com (left-bias). The bias categorization of these media outlets is determined by assessments from MediaBiasFactCheck.com (MBFC). Table 1 organizes corresponding statements from the three articles in rows based on their semantic similarity. If a statement is not mentioned by a particular news source, the corresponding cell in the table is left empty.
By examining Table 1, we can observe instances of cherry-picking. For example, in the first row, all three news outlets mention that President Trump plans to attend the Davos forum in January. However, the third news outlet, cnbc.com (left-bias), provides additional information about Trump skipping the previous year’s gathering and attending the forum in 2020, which is not mentioned by the other outlets. In another example, the second row highlights how the left-biased news outlet, cnbc.com, includes extra information about Trump blaming his no-show on the partial government shutdown, while the other outlets do not mention this detail. Moreover, empty cells in the table indicate that certain statements are not mentioned by reuter.com (center) and newsmax.com (right-bias). This inclusion of information from cnbc.com (left-bias) further contributes to cherry-picking.
reuters.com (Center) newsmax.com (Right-bias) cnbc.com (Left-bias)
U.S. President Donald Trump plans to attend the annual Davos economic forum in January.
U.S. President Donald Trump plans to attend the annual Davos economic forum in January.
After skipping a gathering of the world’s most elite in Davos, Switzerland last January, President Donald Trump will attend the World Economic Forum in 2020.
Trump had to cancel his plan to attend the annual gathering of global economic leaders early this year due to a government shutdown.
Trump had to cancel his plan to attend the annual gathering of global economic leaders early this year due to a government shutdown.
Trump blamed his no-show last time on the partial government shutdown that was triggered by a funding dispute over a proposed wall along the United States’ southern border.
He attended the Davos forum last year.
He attended the Davos forum last year.
The exact date of when Trump would participate was unclear.
The exact date of when Trump would participate was unclear.
A spokesperson for the White House did not immediately respond to a request for comment about the president’s plans to attend next year.
Davos may be one of the few foreign trips that Trump takes in 2020.
Davos may be one of the few foreign trips that Trump takes in 2020.
The World Economic Forum in the Swiss ski resort town of Davos is scheduled to run January 21-24.
The World Economic Forum in the Swiss ski resort town of Davos is scheduled to run January 21-24.
The 50th annual forum, held Jan. 21 - Jan. 24 will focus on “Stakeholders for a Cohesive and Sustainable World.”
Meantime, while the Davos agenda will look to “create bridges to resolve conflicts in global hotspots,” Trump has made jabs at a number of European allies.
While he has said he was just joking about placing tariffs on cars imported from the European Union, he is still threatening to place tariffs on French specialties like champagne and cheese.
Table 1: A juxtaposition of three news stories covering the same event, published by three sources with different political biases. Bias categorization of these media outlets is based on MediaBiasFactCheck.com (MBFC).
The example in Table 1 demonstrates two important insights that guide our approach. color=yellow]Do you think the table indeed demonstrates this point? If so, it might still be helpful to refer to the table more concretely here, even though the details are explained in Appendix. First, cherry-picked statements can be potentially discerned by their importance to the event being covered. Therefore, we exploit statement importance in detecting cherry-picking in news reports. Note that stories often face constraints on space due to factors such as reader’s attention span, which results in prioritizing the most important statements to appear in an article. Second, the importance of a statement is relevant, contextualized by whether it is discussed or not in other stories about the same event—simply put, together all stories depict the full picture. For a comprehensive and representative context, it is crucial to incorporate narratives from unbiased sources or sources that possess a bias different from the source of the given narrative Lee and Lee (1939).
Cluster #1
$\cdot$ President-elect Joe Biden plans to release nearly all available doses of the COVID-19 vaccines after he takes office.
$\cdot$ President-elect Joe Biden plans to release almost all vaccine doses immediately.
$\cdot$ President-elect Joe Biden will aim to release every available dose of the coronavirus vaccine when he takes office.
$\cdot$ Joe Biden will release most available Covid-19 vaccine doses to speed delivery to more people when he takes office.
Cluster #2
$\cdot$ The House voted to override President Trump’s veto of a \$740 billion defense spending and policy bill.
$\cdot$ The House of Representatives on Monday voted to override President Trump’s veto of the National Defense Authorization Act for Fiscal Year 2021.
$\cdot$ House Votes to Override Trump’s Veto of 2021 Defense Policy Bill.
$\cdot$ The House voted late Monday to override President Donald Trump’s veto of a defense spending bill for 2021.
Table 2: Two examples of statement clusters.
We categorize cherry-picking into two primary forms: overemphasis and underemphasis. color=yellow]Several problems with regard to under- vs. over-emphasis. These are worth pondering after paper submission. 1) Our dataset is believed to have very few examples of overempahsis. It is unclear whether it is due to the particular events or particular ways of collecting data. Or whether it is a generally applicable observation. 2) and we didn’t explain what reasons made underemphasis more prevalent, regardless of whether it is due to the data specifically or generally applicable. 3) if it is general, are there studies that agree with our claim? Overemphasis occurs when the coverage of a particular event fixates on non-essential statements and magnifies them, giving the impression that they are significant components of the event Fleming (1995); Lee (1945). Conversely, underemphasis occurs when a story entirely excludes or slants a crucial fact from an event coverage (Sproule, 2001; Lakomy, 2020). The detection of overemphasis requires not only evaluating a statement’s importance but also analyzing its presence in the text, e.g., exaggeration, repetition, and the amount of space within a report the statement occupies. Studying the presence of text requires a different approach than studying its absence. For instance, detecting exaggeration requires computing lexical features which is not applicable to the analysis of textual absence since there is no text to being with. We will address cherry-picking by overemphasis in future work and focus on cherry-picking by underemphasis in this study. color=red]The analysis of overemphasis vs underemphasis is unconvincing. Something present in one article can be missing in another; vice versa. Given their opposite nature, your explanantion doesn’t make a lot of sense to me. To identify underemphasis, analysis of textual absence using context from other narratives and sources is required. To find important statements in an event, one needs to read stories from unbiased sources, then evaluate the importance of each statement within the event based on their understanding of the story they read.
[…]
color=green]Don’t you need to say what one can do using $c_{i}$, like what you did in experiments? It is not the “final output.”. Israa: added
4 Dataset
Figure 1: Cherry-picking data collection interface.
We follow the principles in the previous section in both dataset collection and model training. To train our supervised models to predict the importance of a statement in an event given some context, we collected and curated a novel cherry-picking detection dataset Cherry. Our dataset is composed of 3,346 examples in total. Every example in the dataset contains a statement $s\in S_{e}$ (i.e., a single sentence from a news report of the event $e$), an event context $d\in e$ which contains one or multiple unbiased articles covering the event, and an importance label $Y$ that indicates whether the statement is important to the event or not. To determine which outlets are unbiased, we utilized categorization of media outlet bias based on mutual agreement from three sources—AllSides.com, AdFontesMedia.com, and MBFC, which were discussed in Section 2. color=green]still vague: if MBFC says XYZ.com is unbiased but AllSides.com says it is biased, what do you do?
4.1 Data Collection
To generate the examples in the dataset, we compiled a list of 41 noteworthy news sources, such as CNN.com, Reuters.com, and Foxnews.com, selected for their high publication frequency and distinct political biases as determined by the ratings from the aforementioned three websites.  [^1] The full list of the 41 sources can be found at https://github.com/emnlp-cherry/Cherry. Subsequently, we employed automated tools to gather all news articles from the chosen sources, encompassing the time span between December 2019 and January 2021. Next, we used DBSCAN Ester et al. (1996) to cluster the collected articles into events based cosine similarity between article vectors, where the articles were vectorized by using concatenated BERT Devlin et al. (2019) and TF-IDF representations on the headline and the first paragraph of each article.
After a manual inspection of the generated clusters, we curated a subset of 82 clusters, each corresponding to a distinct controversial event. Examples of such events include the January 6th, 2021 U.S. Capitol attack, the alleged foreign intervention in the 2020 U.S. presidential elections, and the cyberattacks targeting the U.S. Treasury in 2020.  [^2] The full list of the 82 events is available at https://github.com/emnlp-cherry/Cherry.
For each event, we segmented the articles into statements, and then clustered the statements based on semantic similarity. Similar to how articles were clustered, we used BERT and TF-IDF to represent statements and applied DBSCAN to cluster them based on cosine similarity. Table 2 displays a sample of clustered statements within a single event. The statement-clustering functioned as a data augmentation technique, where the collected label for a single statement was cast over all the statements in the same cluster. As a result, multiple examples were labeled using a single example labeling effort.
The data collection, clustering, and segmentation pipeline is depicted in Figure 2. After clustering the statements, we fed them along with their respective context into our custom data annotation tool, as explained in Section 4.2.
4.2 Data Annotation
We developed a data annotation tool  [^3] https://idir.uta.edu/cherrypicking/ to collect labels for statements, indicating their importance to the corresponding events. Upon entering an event ID, annotators are presented with one or two neutral news articles from unbiased sources for the event and a cluster of statements related to the event, as shown in Figure 1. Annotators are required to thoroughly read the article(s) and gain a comprehensive understanding of the event covered. The instruction provides guidance on categorizing the statement cluster on the right-hand side of Figure 1. Annotators are instructed to assign one of the five labels to the statements: (1) very important, (2) kind of important, (3) not very important, (4) the excerpts might be incorrect, and (5) I am not sure. If the statement cluster is unrelated to the event, lacks meaningfulness, or contains facts with conflicting meanings, annotators should label it as “the excerpts might be incorrect.” Annotators can select “I am not sure” if they feel uncertain about the importance of the statements. After selecting a label, annotators can proceed by clicking the “Submit and go to next example” button to work on the subsequent statement clusters associated with the same event. The annotators are one faculty and nine students from our research lab.
Figure 2: Data collection and preparation pipeline.
Following the collection of labels from annotators, we filtered out examples whose agreement ratio was less than 75% or whose number of annotators was less than 3. Agreement ratio is given by the number of majority votes divided by the total number of voters.
4.3 Dataset Statistics
Considering the inherent difficulty of the problem, we opted to explore various binary and multinomial classification configurations as depicted in Table 3. This involved merging certain labels into a single class or excluding specific labels from the dataset. The class distribution for each of the four classification configurations is provided in Table 4. For each configuration, we split the dataset randomly into 85% for training and 15% for testing.
The dataset contains couples consisting of statements and contexts from multiple events. Consequently, if multiple statements stem from the same event, the context is repeated within the dataset. This means that many examples in the dataset share the same context, which comprises half of the example. Hence, we split the dataset between testing and training by event not by stratification to avoid data leakage.
Conf. Class 1 Class 2 Class 3
1
{(1)}
{(2),(3),(4),(5)}
2
{(1)}
{(2),(3)}
3
{(1)}
{(2),(3)}
{(4),(5)}
4
{(1)}
{(2)}
{(3)}
Table 3: Combinations of labels used in different classification configurations. (1) very important, (2) kind of important, (3) not very important, (4) the excerpts might be incorrect, (5) I am not sure.
Conf. Class 1 Class 2 Class 3
1
2175 (64%)
1232 (36%)
2
2175 (65%)
1171 (35%)
3
2175 (64%)
1171 (34%)
61 (2%)
4
2175 (65%)
667 (20%)
504 (15%)
Table 4: Class distribution for each of the four classification configurations.
5 Models
To automatically detect cherry-picking, our models emulate the approach taken by human annotators by reading a neutral context about the event then assessing a given statement’s importance considered that context. To achieve that, we trained two language models, BERT Devlin et al. (2019) and Longformer Beltagy et al. (2020), with a sentence pair classification setting on our dataset, as shown in Figure 3. The sequence pair consists of the the statement text, and the context text. If the token limit is exceeded, the sequence is right-truncated. The encoded sequence is represented as a vector obtained from the output of the [CLS] token. This vector is subsequently passed through a linear layer to produce a class probability.
Longformer, unlike BERT, is a language model that exhibits linear scaling with the length of the input sequence. As a result, it enables the processing of lengthier documents, accommodating approximately 4096 tokens compared to BERT’s limit of 512 tokens. The primary factor that distinguishes Longformer as a computationally efficient variant from BERT is its novel self-attention mechanism. Instead of applying self-attention on every token in the input sequence, Longformer saves memory and processing time by applying self-attention over a sliding window or dilated window of tokens. It relies on the upper encoder layers to capture the full context from all the self-attention windows. Additionally, to learn task-specific representations, Longformer uses global attention on preselected tokens form the input sequence. These tokens attend to every other token in the sequence, and every token in the sequence attends only to these preselected tokens.
Figure 3: Automatic cherry-picking detection using sentence pair classification architecture.
Longformer proves to be more advantageous for our requirements compared to BERT. This is primarily due to Longformer’s capability to handle lengthier texts, which aligns well with our typically long sequences of context (i.e., complete articles). Additionally, Longformer’s global attention mechanism facilitates the learning of task-specific representations Beltagy et al. (2020), further enhancing its suitability for our needs. color=green]This paragraph was not very clear to me. Will check later. Israa: rewritten
By training these language models, we can assess the importance of a given statement within an event. The cherry-picking detection pipeline identifies cherry-picked statements $c_{i}$ within a document $d_{i}$ given a list of events $E$ that have been clustered using semantic similarity based on:
$c_{i}=\{\exists s\in S_{e}:f(s,d)\geq T\}-d_{i}$ (1)
where $f$ is one of our supervised models, and $T$ is a threshold on the importance score. Events are obtained from a 24-hour snapshot of published news articles from a specified set of sources. The pipeline finds $c_{i}$ by first creating the universal set of segments $S_{e}$ by segmenting all documents in event $e$. Then each segment in $S_{e}$ is scored for importance using the trained supervised models and the context $d$ collected from a neutral article in the event. Finally, each important segment is verified for its presence in each article using semantic similarity. If the segment is absent in an article, it is appended to the list of cherry-picked statements associated with that article.
6 Experimentation and Results
Importance granularity.
Recognizing the challenge associated with manually assessing the importance of a statement within an event, we anticipate that the performance of models trained on varying levels of statement importance granularity will differ as well. For instance, we expect supervised models to be more precise in drawing a boundary between the labels very important versus everything else compared to drawing a boundary between more fine-grained importance labels such as very important versus kind of important. To examine our models’ capabilities across different levels of importance granularity, we trained our top-performing model (i.e., Longformer-Large) using the four classification configurations outlined in Table 3. Results in Table 5 show that the difficulty of learning a decision boundary between the different labels decreases when we go less granular and merge more classes, especially kind of important and not very important classes. This is attributed to the high fuzziness of these two classes. Therefore, we utilize configuration 1 as indicated in Table 3 in all subsequent experiments.
Conf. # Acc. F-1
1 0.897 0.887
2 0.853 0.828
3 0.875 0.898
4 0.728 0.717
Table 5: Performance of the Longformer model in the four classification configurations.
The contribution of context in predicting a statement’s importance.
To evaluate the contribution of context to the performance of the models in predicting a statement’s importance, we employed a BERT-base model as our baseline, with the fact alone (i.e. no context) as its input sequence. Then we created another variant of the same model by adding the context to the input sequence and separating the sequence by the [SEP] token. Both variants of the model allow for a sequence length of 128 tokens. Our hypothesis is that just as humans can make more accurate judgments regarding the importance and cherry-picking of a statement when provided with an unbiased context Lee and Lee (1939), supervised models are anticipated to exhibit similar behavior. Results in the first two rows of Table 6 show a 28% increase in performance in terms of F-1 score when introducing context to the model through the input sequence. [^4] All the code base, experiments and results in this section are available on: https://github.com/emnlp-cherry/Cherry/.
The effect of context length in predicting a statements importance.
To understand the effect of longer context to the model’s ability to predict the importance of a statement, we conducted experiments with varying context lengths. Our assumption is that since the most important information is typically found at the beginning of an article, increasing context length up to a certain point should improve the model’s ability to predict a statement’s importance. To implement this experiment, we truncated the sequence from the context side only, and set the maximum sequence length at different values. For all BERT models we used the BERT-base variant. We fixed the learning rate at 2e-05 and batch size at 8 for all BERT and Longformer variants. Results in Table 6 show an improvement in performance in terms of accuracy and F-1 score as the context size increases. However, this improvement stops once all important information are exhausted as shown in the seventh row of Table 6. This is because the model starts to see less important information which increases model’s confusion. Another take away from this experiment is that we can further improve performance when using Longformer architecture and this is attributed to its global attention mechanism that captures better representations for the long context sequences. This is shown in rows 3 to 6 of Table 6 where both BERT and Longformer were tested using the same context length. Furthermore, results in the last row of Table 6 show there is still a room for a significant improvement in performance when using larger language models.
Model Seq. length Acc. F-1
BERT-Base 128,no context 0.619 0.617
BERT-Base 128 0.789 0.792
BERT-Base 256 0.826 0.811
BERT-Base 512 0.836 0.840
Longf.-Base 256 0.832 0.836
Longf.-Base 512 0.843 0.846
Longf.-Base 1024 0.813 0.819
Longf.-Large 512 0.897 0.887
Table 6: Performance of the BERT and Longformer models with varying context lengths.
Global attention locations.
To evaluate the influence of the global attention mechanism on predicting a statement’s importance, we tested our best model with all possible combinations of global attention locations within the input sequence (i.e., the classification token, the fact tokens, and the context tokens). We maintained the sequence length of 512 but decreased the batch size to 4 in this experiment to ensure memory can handle applying global attention on all the sequence tokens. Results in Table 7 show that we can still maintain the same performance but optimize memory and inference time Beltagy et al. (2020) by applying global attention on the statement’s and classification tokens only. For this reason, we fix global attention to these two locations in all other experiments. Moreover, we can optimize further and trade off more efficiency at the expense of a slight decrease in performance by applying global attention on the classification token only.
Global attention locations Acc. F-1
[CLS] 0.819 0.820
[CLS] + statement 0.831 0.837
[CLS] + context 0.683 0.680
[CLS] + statement + context 0.837 0.838
Table 7: Performance of the Longformer model using different global attention locations.
End-to-end cherry-picking detection.
To assess the performance of the complete cherry-picking detection pipeline, we examined the relationship between two variables, $x$: Cherry-picking detected by our automatic pipeline in each news outlet and $y$: the bias scores of the outlet collected from MBFC and AllSides. Our assumption is that since cherry-picking is one form of bias, there should be some level of correlation between the two variables. However, no strong correlation between the two variables is anticipated. This is because bias scores from MBFC and AllSides are estimated by measuring all other different types of bias indicators such as lexical cues present in the text itself. Additionally, MBFC and AllSides update their bias scores regularly based on new samples of articles, while our cherry-picking scores are based on fixed data we collected in the period Dec. 2019 to Jan 2021. Hence, there is a mismatch between the data analyzed to generate bias scores and cherry-picking scores, which can be addressed in future work.
To test the relationship between variables $x$ and $y$, we ran the full cherry-picking detection pipeline described in formula 1 employing our top-performing supervised model on 2,453 events comprised of about 97,000 statements collectively. We ensured that these events did not overlap with the Cherry dataset utilized for training our supervised models. Next, we calculated the average number of cherry-picked statements for each outlet across all events, resulting in a final score for the variable $x$. Finally, since variable $y$ is ordinal, we calculated the Spearman’s correlation coefficient $r$ for variables $x$ and $y$. Results in Table 8 show a positive moderate correlation approaching statistical significance between our cherry-picking scores and bias scores from AllSides.com. Additionally, there is a positive weak quasi-significant correlation between our cherry-picking scores and bias scores from MBFC.
Lastly, we hypothesize that the average cherry-picking score for biased outlets, belonging to a particular bias band, should align with their bias intensity. For instance, we anticipated observing a higher average cherry-picking score among outlets categorized as “left bias” compared to those classified as “left-center bias.” To test that, we calculated the average and standard deviation of the cherry-picking score of all outlets under each bias category as Table 9 shows. Results show some increase in cherry-picking scores that aligns with the intensity of bias in the analyzed outlets. Nevertheless, the observed pattern is not as sharp as anticipated due to the limited sample sizes, highlighting the need for future analysis involving larger sample sizes.
$y$ r P-value
MBFC 0.28 0.10
AllSides 0.32 0.06
Table 8: Correlation results between $x$: cherry-picking detection pipeline scores and $y$: bias scores from MBFC and AllSides.com.
Bias category Mean STD Sample size
Left 15.12 4.39 6
Left center 10.32 3.05 15
Right 8.91 4.09 7
Right center 8.33 1.21 5
Center 8.44 0.30 2
Table 9: The mean and standard deviation of cherry-picking scores aggregated by bias category.
7 Conclusions and Future Work
Cherry-picking statements in news articles is challenging to spot manually, especially due to the need for careful examination of other stories discussing the same event. This study introduces a novel approach to automate the detection of cherry-picking in news reports. Our approach focuses on comparing multiple news articles that pertain to the same event and identifying the omission of crucial statements. To facilitate our research, we have constructed the first cherry-picking detection dataset. The preliminary results of our models demonstrate promising outcomes in detecting cherry-picked statements. Currently, our work specifically addresses the detection of cherry-picking in news articles based on underemphsis of statements.
In future endeavors, we aim to expand this work by also considering overemphasis, where news reports intentionally include unimportant or irrelevant statements regarding an event. Furthermore, we plan to leverage generative language models to further improve the detection of cherry-picking by training the models to generate alternative narratives that are devoid of cherry-picked content. By making these enhancements, we strive to develop a more comprehensive and effective methodology for detecting cherry-picking in news articles.
Limitations
Dependence on semantic similarity thresholds and adjustable parameters. A notable limitation of the end-to-end cherry-picking pipeline, as outlined in Section 5, is its dependence on semantic similarity thresholds and other adjustable parameters. This can be observed in the application of clustering techniques during the creation of events from a set of articles. If a news outlet covers an event through multiple articles and one of these articles is not clustered into the correct event, the pipeline may overlook important information from the coverage of that outlet.
Semantic similarity threshold Moreover, when determining whether an article mentions a specific statement, if the threshold for semantic similarity is not well-tuned, similar statements in the article may not be detected and consequently be considered cherry-picked.
Uncertainty of the supervised model. The supervised model’s uncertainty should be calibrated before relying on the importance threshold value in formula 1. This ensures that only statements deemed important by the model with a high level of confidence are considered when assessing cherry-picking in a given news report.
Definition and tokenization of statements. To ensure the precise identification of cherry-picked statements, it is essential to have a clear definition and an appropriate tokenization method for segmenting text into statements. Currently we consider every sentence as a statement, but this approach may not always be accurate since multiple sentences can collectively form a single statement.
Reliance on a single unbiased source for event context. Our approach primarily relies on Reuters.com as the only unbiased source from which we derive event context. However, collecting context from multiple sources does not require any change in the proposed methodology.color=green]need to break this long paragraph and itemize the individual limitations. Also, make sure to proofread it. I corrected some typos and grammar errors; there must be more.
Ethics Statement
While our work focuses on detecting bias by omission, it inevitably incorporates a human element in two stages. First, At the data collection stage, human coders are asked to indicate whether a statement is important to the event. This may result in a dataset that contains the inherent subjectivity and personal perspectives of the human annotators themselves. We are aware that annotators’ implicit biases can influence the annotations they provide. Therefore, we tried to alleviate this influence by hiding outlets names from segments in the annotation tasks, in addition to providing annotators with a neutral context to read before they assign labels. Second, bias categorizations we use from MBFC, AllSides, and AdFontesMedia rely on human annotators as well. Hence, the bias of these sources can be inherently embedded into our dataset. While we hold the belief that the sources we rely on exhibit high quality, it is important to recognize that their reliability and credibility may decline due to various unforeseen factors.